OpenStudy (anonymous):

Use the quadratic formula to solve the equation. 2x^2 − 7x − 6 = 0 x= (smaller value) x= (larger value)

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

x=-b+and-sqrtb^2-4ac a=2 b=7 c=6

OpenStudy (anonymous):

over 2a

OpenStudy (anonymous):

x=-bmoreorless sqrtb^2-4ac over 2a

OpenStudy (anonymous):

You have the wrong b and c. You forgot your negatives: a=2, b=-7, c = -6 Therefore \[x=\frac{-(-7)\pm \sqrt{7^2 - 4(2)(-6)}}{2(2)}\]

OpenStudy (anonymous):

\[2x^2 -7x-6=0\] first determine the values of a, b, and c. a= 2, b=-7 and c=-6 and substitute them to the quadratic formula: \[x= (-b \pm \sqrt{b^2 -4ac}) \div2a \] hence, \[x= {7 \pm \sqrt{49-4(2)(-6)}} \div (2*2) \] \[x=(7 \pm \sqrt{97}) \div 4\]

OpenStudy (anonymous):

\[\left\{x\to \frac{1}{4} \left(7-\sqrt{97}\right),x\to \frac{1}{4} \left(7+\sqrt{97}\right)\right\}=\{x\to -0.712214,x\to 4.21221\} \]

OpenStudy (anonymous):

x=4.212 is right but x=0.712214 its wrong

OpenStudy (anonymous):

Replacing x in the expression:\[2x^2-7x-6 \]with\[\frac{1}{4} \left(7-\sqrt{97}\right) \]gives\[\frac{1}{8} \left(7-\sqrt{97}\right)^2-\frac{7}{4} \left(7-\sqrt{97}\right)-6 \]Expand the first term and simplify.\[\frac{73}{4}-\frac{7 \sqrt{97}}{4}-\frac{7}{4} \left(7-\sqrt{97}\right)-6 \]Expand the second term and combine with the third\[\frac{73}{4}-\frac{7 \sqrt{97}}{4}\left(-\frac{73}{4}+\frac{7 \sqrt{97}}{4}\right) \]The sum of the remaining terms is zero.